### v1

# Determinant Estimation under Memory Constraints and Neural Scaling Laws

Abstract
Calculating or accurately estimating log-determinants of large positive semi-definite matrices is of fundamental importance in many machine learning tasks. While its cubic computational complexity can already be prohibitive, in modern applications, even storing the matrices themselves can pose a memory bottleneck. To address this, we derive a novel hierarchical algorithm based on block-wise computation of the LDL decomposition for large-scale log-determinant calculation in memory-constrained settings. In extreme cases where matrices are highly ill-conditioned, accurately computing the full matrix itself may be infeasible. This is particularly relevant when considering kernel matrices at scale, including the empirical Neural Tangent Kernel (NTK) of neural networks trained on large datasets. Under the assumption of neural scaling laws in the test error, we show that the ratio of pseudo-determinants satisfies a power-law relationship, allowing us to derive corresponding scaling laws. This enables accurate estimation of NTK log-determinants from a tiny fraction of the full dataset; in our experiments, this results in a $\sim$100,000$\times$ speedup with improved accuracy over competing approximations. Using these techniques, we successfully estimate log-determinants for dense matrices of extreme sizes, which were previously deemed intractable and inaccessible due to their enormous scale and computational demands.
1 Introduction
Many quantities of interest in machine learning require accurate estimation of the log-(pseudo-)determinant of large dense positive (semi-)definite matrices, often indexed by some number of datapoints. These quantities arise in a number of tasks, including training Gaussian processes (GPs) and other kernel-based methods (Wang et al., , 2019), graphical models (Rue & Held, , 2005), determinantal point processes (Kulesza et al., , 2012), and model comparison and selection techniques (Hodgkinson et al., 2023b, ). Problems where the calculation cannot be avoided can often be reduced to computing a volume form, which is the case for tasks in statistical mechanics (Mézard & Montanari, , 2009) or Bayesian computation (Gelman et al., , 2013), or applications of the Karlin-McGregor theorem (Karlin & McGregor, , 1959) including determinantal point processes. Similarly, when training GPs via empirical Bayes (Rasmussen & Williams, , 2006), the log-determinant term is the most difficult to compute. In many applications, these matrices are not only dense but highly ill-conditioned. This renders methods that leverage sparsity inappropriate and makes accurate estimation of small eigenvalues crucial, since errors in these small values are magnified at log-scale. Much of the research focus has concentrated on approximation techniques that ameliorate the cubic computational complexity of these log-determinant calculations. Methods leveraging stochastic expansions that rely upon matrix-vector multiplications—such as Lanczos-based methods (Ubaru et al., , 2017)—have been particularly successful, yielding approximate methods that scale linearly (Dong et al., , 2017), and methods leveraging Taylor’s expansions (Fitzsimons et al., , 2017) have also been proposed. However, accuracy can suffer when they encounter common ill-conditioning pathologies. Crucially, time complexity isn’t the only bottleneck at play: large-scale matrices found in modern applications often hit the memory wall before computational cost becomes an issue (Gholami et al., , 2024).
Recently, the empirical Neural Tangent Kernel (NTK) has become a prominent theoretical and practical tool for studying the behavior of neural networks both during training and at inference (Novak et al., , 2022). The Gram matrix associated with the NTK has been used in lazy training (Chizat et al., , 2019) and shows promise as a tool to obtain uncertainty quantification estimates (Immer et al., , 2021). Its log-determinant—highly sensitive to small eigenvalues—has also received recent attention, both as a quantity of interest in model selection techniques that rely upon marginal likelihood approximation (Immer et al., , 2023; Hodgkinson et al., 2023b, ) and as a way to quantify the complexity of a learning problem (Vakili et al., , 2022). Similarly, both the NTK log-determinant and closely related quantities have recently appeared in quantification of generalization error via PAC-Bayes bounds (Hodgkinson et al., 2023c, ; Kim et al., , 2023). While several approximations have been proposed (Mohamadi et al., , 2023), convergence has only been shown in spectral norm, so they can only be expected to capture the overall behavior of the full NTK.
In practice, computing the log-determinant of the empirical NTK’s Gram matrix is a formidable task: besides suffering from the pathologies that ill-conditioned matrices are subject to at scale, the NTK appears to have its own peculiarities that make the problem particularly challenging. Indeed, the task is considered so impenetrable as to be universally avoided, since the NTK corresponding to most relevant datasets cannot be stored in memory. For example, storing the NTK corresponding to the relatively small MNIST dataset requires 2.9 terabytes of memory. The equivalent object for ImageNet-1k requires 13.1 exabytes, an order of magnitude larger than CERN’s current data storage capacity (Smith, , 2023). While it is appealing to consider only a small subset of the training dataset when considering such approximations, these estimates for the log-determinant can incur significant bias. Furthermore, we will find that leading approaches for dealing with log-determinants of empirical NTK Gram matrices using sketching and other Monte Carlo approximations tend to be highly inaccurate.
Contributions.
The central contributions of this work address the issues faced when computing log-determinants of large matrices. We are primarily interested in matrices that cannot fit into memory. To compute an accurate baseline, we derive MEMDET, a memory-constrained algorithm for determinant computation. This facilitates exact calculation of log-determinants of NTK Gram matrices corresponding to neural networks with several million parameters, for the first time on datasets this large. We then provide a novel approximation technique based on the scaling behavior of a wide class of kernel matrices, by appealing to neural scaling laws. Scaling laws have played a prominent role in machine learning theory and practice, providing insight into the asymptotic behavior of generalization error (Li et al., , 2023; Vakili et al., , 2022), as well as guidance for compute-optimal resource allocation when deploying deep learning models at scale (Kaplan et al., , 2020; Hoffmann et al., , 2022). Specifically,
  • we derive a hierarchical memory-constrained algorithm for large-scale computation of log-determinants, which we name MEMDET;
  • under mild assumptions, we derive scaling laws for the ratio of pseudo-determinants of kernel matrices containing different subsets of the same dataset, enabling both a corresponding law of large numbers and central limit theorem for normalized log-determinants;
  • leveraging these scaling laws, we propose FLODANCE, a novel algorithm for accurate extrapolation of the log-determinant from a small fraction of the full dataset;
  • we demonstrate the practical utility of our method by approximating the NTK corresponding to common deep learning models; and
  • we provide a high-performance Python package detkit [^1] For details see Appendix I., which implements the presented algorithms and can be used to reproduce the results of this paper.
Crucially, we demonstrate that our approximation technique is able to obtain estimates with lower error than incurred by reducing the numerical precision used in explicit computation. Our approximation techniques render an impractical task virtually routine, as shown in Table 1. To the best of our knowledge, this is the first time that the full empirical NTK corresponding to a dataset of this scale has been computed.
Table 1: Comparison of stochastic Lanczos quadrature (SLQ, with degree $l$, $s$ Monte Carlo samples, and full re-orthogonalization), MEMDET (Algorithm 3), and FLODANCE (Algorithm 1) on a dense $500000$\times$500000$$ NTK matrix for a ResNet50 model trained on CIFAR-10 with $50000$ datapoints. MEMDET computes the exact log-determinant and serves as the benchmark, with relative errors of other methods measured against it. Costs and wall time are based on an NVIDIA H100 GPU (\$2/hour) and an 8-core 3.6GHz CPU (\$0.2/hour) using Amazon pricing.
Method Rel. Error Est. Cost Wall Time
Name Settings TFLOPs
SLQ $l=100$, $s=104$ 5203 55% \$83 1.8 days
MEMDET LDL, $n_{b}=32$ 41,667 0% \$364 8.8 days
FLODANCE $n_{s}=500,q=0$ 0.04 4% \$0.04 1 min
FLODANCE $n_{s}=5000,q=4$ 41.7 0.02% \$4 1.5 hr
The remainder of this document is structured as follows. In Section 2, we discuss the computational issues faced when computing determinants at scale, and we derive our MEMDET algorithm based on block LU computation of the log-determinant. Section 3 contains the appropriate scaling laws for pseudo-determinants of interest, the corresponding LLN and CLT for their logarithms, and the FLODANCE algorithm for their approximation. Numerical experiments are presented in Section 4, and we conclude in Section 5.
2 Computing Determinants at Scale
The computation of determinants of large matrices, particularly those expected to be highly ill-conditioned, is widely considered to be a computationally “ugly” problem, which should be avoided, if possible (Axler, , 1995). However, this is not always an option, for example when the determinant represents a volume form, is required in a determinantal point process, or plays a role in training GPs via empirical Bayes. Although certain limiting behaviors under expectation are well characterized (Hodgkinson et al., 2023a, ), they provide only a coarse approximation in finite-sample settings. The quadratic memory cost and cubic computational complexity of naïve implementations make exact computation prohibitive, necessitating the use of lower-precision or randomized methods at scale. While approximate methods can be useful in their own right, exact computation remains essential, at the very least, to establish meaningful baselines. In our experiments, due to the pathological spectral behavior of the matrices we consider, we will see that our proposed exact method is comparable in speed to the state-of-the-art Monte Carlo approximation techniques.
2.1 Low-Precision Arithmetic
Among the most common techniques for circumventing memory and computational bottlenecks when dealing with with large-scale calculations in numerical linear algebra is to cast numerical values into a lower precision, usually 32-bit, 16-bit, or even 8-bit floating point values, instead of 64-bit (double precision) values that would otherwise be used.
Computations of the log-determinant in mixed precision do not generally incur significant error. However, in our setting, the matrix of interest is realized as the product $\mathbf{J}\mathbf{J}^{\intercal}$, where $\mathbf{J}$ is the Jacobian. The formation of the quadratic is a well-known source of approximation error, and should be avoided if possible: paraphrasing Higham, (2022), if $\delta$ is the round-off used in choice of floating point arithmetic, then for $0<\varepsilon<\sqrt{\delta}$ we can consider the simple case of $\mathbf{J}=\left[\begin{array}[]{cc}1&\varepsilon\\ 1&0\end{array}\right]$, we have
$\mathbf{J}\mathbf{J}^{\intercal}=\left[\begin{array}[]{cc}1+\varepsilon^{2}&1\\ 1&1\end{array}\right],\quad\mathsf{fl}(\mathbf{J}\mathbf{J}^{\intercal})=\left[\begin{array}[]{cc}1&1\\ 1&1\end{array}\right],$
where $\mathbf{J}\mathbf{J}^{\intercal}$ rounds to the singular $\mathsf{fl}(\mathbf{J}\mathbf{J}^{\intercal})$. In general, significant precision loss should occur if $\mathsf{cond}(\mathbf{J}\mathbf{J}^{\intercal})>\delta^{-1}$, where $\delta\approx 10^{-3}$, $10^{-6}$, and $10^{-15}$ for 16-bit, 32-bit, and 64-bit precisions respectively. In the Gram matrices we consider, $\mathsf{cond}(\mathbf{J}\mathbf{J}^{\intercal})>10^{12}$. Due to the scale of the problems of interest here, the sheer size of the Jacobian (requiring at least hundreds of terabytes of space) makes directly operating on $\mathbf{J}$ impractical. This necessitates breaking conventional wisdom and explicitly forming the Gram matrix. We will see later the incurred cost to accuracy when working in low precision.
2.2 MEMDET: A Memory-Constrained Algorithm for Log-Determinant Computation
Conventional methods for computing the determinant of matrices include LU decomposition (for generic matrices), LDL decomposition (for symmetric matrices), and Cholesky decomposition (for symmetric positive-definite matrices). These methods also simultaneously provide the determinants of all sub-matrices $\mathbf{M}_{[:k,:k]}$, consisting of the first $k$ rows and columns ($k=1,\dots,m$) of the $m\times m$ matrix $\mathbf{M}$. This will be beneficial in our applications.
The LU decomposition has a computational complexity of approximately $\frac{2}{3}m^{3}$, while both LDL and Cholesky decompositions have a complexity of approximately $\frac{1}{3}m^{3}$. While computational complexity is a primary concern for large-scale determinant computation, memory limitation poses an even greater challenge. These methods require substantial memory allocation, either the size of the array if written in-place, or twice the size if the input array is preserved.
To address this, we present MEMDET, a memory-constrained algorithm for large-scale determinant computation. Below is a sketch of the algorithm, with details of its implementation, memory and computational aspects provided in Appendix C, and a description of the software implementing this algorithm provided in Appendix I. We focus on the algorithm for generic matrices using LU decomposition, though the LDL and Cholesky decompositions follow similarly.
Consider the $2\times 2$ block LU decomposition of $\mathbf{M}$ (see for instance Dongarra et al., 1998b (, Chapter 5.4)) with the first block, $\mathbf{M}_{11}$, of size $b\times b$, $b<m$, as
$\mathbf{M}=\begin{bmatrix}\mathbf{M}_{11}&\mathbf{M}_{12}\\ \mathbf{M}_{21}&\mathbf{M}_{22}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{11}&\mathbf{0}\\ \mathbf{L}_{21}&\mathbf{I}\end{bmatrix}\begin{bmatrix}\mathbf{U}_{11}&\mathbf{U}_{12}\\ \mathbf{0}&\mathbf{S}\end{bmatrix},$ (1)
where $\mathbf{L}_{11}$ is lower triangular, $\mathbf{U}_{11}$ is upper-triangular, and $\mathbf{I}$ is the identity matrix. The blocks of the decomposition are obtained by computing $\mathbf{M}_{11}=\mathbf{L}_{11}\mathbf{U}_{11}$, solving lower-triangular system $\mathbf{U}_{12}=\mathbf{L}_{11}^{-1}\mathbf{M}_{12}$ and upper-triangular system $\mathbf{L}_{21}=\mathbf{M}_{21}\mathbf{U}_{11}^{-1}$, and forming the Schur complement $\mathbf{S}\coloneqq\mathbf{M}_{22}-\mathbf{L}_{21}\mathbf{U}_{12}$. This procedure is then repeated on the $(m-b)\times(m-b)$ matrix $\mathbf{S}$, treated as the new $\mathbf{M}$, leading to a new $b\times b$ upper-triangular matrix $\mathbf{U}_{11}$ and a smaller Schur complement $\mathbf{S}$ at each iteration. This hierarchical procedure continues until the remaining $\mathbf{S}$ is of size $b$ or less, at which point its LU decomposition is computed. The log-determinant of the entire matrix is the sum of the log-determinants of all $\mathbf{U}_{11}$ blocks at each iteration.
[…]
Figure 1: Schematic diagram of MEMDET illustrating the efficient block processing order for LU decomposition (Algorithm 2, left panel) and LDL/Cholesky decompositions (Algorithm 3, Algorithm 4, right panel). The detailed ordering strategy is described in Appendix D.
To make this memory-efficient, the algorithm is modified to hold only small chunks of the matrix in memory, storing intermediate computations on disk. Suppose $\mathbf{M}$ consists of $n_{b}\times n_{b}$ blocks $\mathbf{M}_{ij}$ of size $b\times b$ where $i,j=1,\dots,n_{b}$. We pre-allocate four $b\times b$ matrices $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, and $\mathbf{S}$ in memory. The $k$-th stage begins by loading $\mathbf{A}\leftarrow\mathbf{M}_{kk}$ from disk and performing an in-place LU decomposition $\mathbf{A}=\mathbf{L}\mathbf{U}$, with $\mathbf{L}$ and $\mathbf{U}$ stored in $\mathbf{A}$. We then compute the Schur complement for all inner blocks $\mathbf{M}_{ij}$, $i,j=k+1,\dots,n_{b}$, by loading $\mathbf{B}\leftarrow\mathbf{M}_{kj}$ and $\mathbf{C}\leftarrow\mathbf{M}_{ik}$ from disk, solving $\mathbf{B}\leftarrow\mathbf{L}^{-1}\mathbf{B}$ and $\mathbf{C}\leftarrow\mathbf{C}\mathbf{U}^{-1}$ in-place, loading $\mathbf{S}\leftarrow\mathbf{M}_{ij}$, and computing $\mathbf{S}\leftarrow\mathbf{S}-\mathbf{C}\mathbf{B}$. The updated $\mathbf{S}$ is then stored back to disk $\mathbf{M}_{ij}\leftarrow\mathbf{S}$, either by overwriting a block of the original matrix, or avoids this by writing to a separate scratchpad space. In the latter case, a cache table tracks whether a block $\mathbf{M}_{ij}$ should be loaded from the original matrix or from the scratchpad in future calls.
The computational cost of this procedure remains the same, independent of the number of blocks, $n_{b}$ (see Section E.1). However, increasing the number of blocks reduces memory usage at the expense of higher data transfer between disk and memory. Efficient implementation minimizes this by processing blocks in an order that reduces the loading of $\mathbf{B}$ and $\mathbf{C}$. The left panel of Figure 1 shows one such order, illustrating the procedure at the $k$-th iteration. Processing the blocks $\mathbf{M}_{ij}$ starts from the last row of the matrix and moves upward. During the horizontal and vertical traverses on the path shown in the figure, the memory blocks $\mathbf{C}$ and $\mathbf{B}$ can remain in memory, avoiding unnecessary reloading. Once the procedure reaches the block $(k+1,k+1)$, the memory block $\mathbf{S}$ can be directly read into $\mathbf{A}$ instead of being stored on disk as $\mathbf{M}_{k+1,k+1}$, initiating the block $\mathbf{A}$ for the next iteration. Consequently, only the blocks shown in blue need to be stored in the scratchpad space, while those in dark blue are already stored in the current state of the algorithm. This algorithm can be further modified to eliminate the need for $\mathbf{A}$ and use the memory space of $\mathbf{S}$ instead, but this would require storing the additional gray blocks on the scratchpad space in addition to the blue blocks. A pseudo code and further efficient implementation details of the presented method can be found as Algorithm 2 in Appendix D.
The algorithms for LDL and Cholesky decompositions follow a similar procedure with necessary adjustments for symmetric and symmetric positive-definite matrices, as detailed in Algorithm 3 and Algorithm 4. These modifications include processing only the lower (or upper) triangular part of the matrix and handling permutations and diagonal scaling in the LDL decomposition. The right panel of Figure 1 shows one possible block ordering for the algorithm for LDL and Cholesky decomposition: the ordering is optimal in minimizing the number of reads and writes, but is not unique. Other block processing orderings with the same amount of data access also exist, and these will be further discussed in Appendix D. A detailed analysis of the computational complexity and memory/disk data transfers of the algorithm is provided in Appendix E. Implementation of MEMDET algorithm can be found in LABEL:list:detkit1.
Although the Neural Tangent Kernel (NTK) matrices we work with are symmetric and positive definite (SPD), we do not solely rely on Cholesky decomposition, despite it being the most suitable method for SPD matrices. This is because NTK matrices can lose their positive-definiteness with even the smallest perturbations causing small eigenvalues to become negative, such as when converting from 64-bit to 32-bit precision for efficient computation. As a result, the Cholesky decomposition becomes unstable and fails, necessitating the use of LU and LDL decompositions, suitable for more general matrices. We note that the block computations of LU decomposition can become unstable as the matrices deviate from symmetric and positive-definiteness (Demmel et al., , 1995), requiring pivoting of the blocks. However, in our empirical study of NTK matrices, we found that LU decomposition works well without block pivoting, though we do consider pivoting within each block.
3 Scaling Law for the Determinant
We now turn our attention to the estimation of the log-determinant of Gram matrices corresponding to covariance kernels. Our primary motivating example is the NTK, an architecture-specific kernel associated with deep neural network models. The connections between neural networks and GP kernels are well-known, particularly the now classical results that networks at initialization induce a GP kernel in the large-width limit, enabling Bayesian inference for these infinite-width networks (Neal, , 1996; Lee et al., , 2018). Similarly, linearization of the gradient-flow dynamics during the late stages of training leads to the derivation of the NTK, whose infinite-width analogue can be shown to be constant during this training phase (Jacot et al., , 2018; Yang, , 2020; Yang & Littwin, , 2021).
The NTK was first derived in the context of neural networks. However, the quantity is well defined for more a general class of functions. For a continuously differentiable function $f_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$, we define its (empirical) NTK to be
$\kappa_{\theta}(x,x^{\prime})\coloneqq\mathbf{J}_{\theta}\big{(}f_{\theta}(x)\big{)}\mathbf{J}_{\theta}\big{(}f_{\theta}(x^{\prime})\big{)}^{\intercal},$ (2)
where $\mathbf{J}_{\theta}(f_{\theta}(x))\in\mathbb{R}^{d\times p}$ is the Jacobian of the function $f_{\theta}$ with respect to the flattened vector of its parameters $\theta\in\mathbb{R}^{p}$, evaluated at the point $x$. Note that the assumption of continuous differentiability is often relaxed in practice, and we have $\kappa_{\theta}(x,x^{\prime})\in\mathbb{R}^{d\times d}$. When computing the NTK across $n_{1}$ and $n_{2}$ datapoints, it is typical to arrange the $n_{1}n_{2}$ matrices as an $n_{1}d\times n_{2}d$ matrix. The scaling laws we derive in the sequel hold for a class of kernel functions satisfying certain decay properties on the eigenvalues of an associated integral operator (see Appendix F for a precise statement, taken from Li et al., (2023)). We remark that this class of kernels includes NTKs, as demonstrated in Bietti & Mairal, (2019); Bietti & Bach, (2021); Lai et al., (2023).
3.1 Neural Tangent Kernels and Scaling Laws
Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a positive-definite kernel. For a sequence of inputs $\{x_{i}\}_{i=1}^{\infty}\subset\mathcal{X}$ and for each $n=1,2,\dots$, let $\mathbf{K}_{n}\coloneqq(\kappa(x_{i},x_{j}))_{i,j=1}^{n}$ be the corresponding Gram matrix for the first $n$ inputs. Our objective is to estimate $\mathsf{logdet}(\mathbf{K}_{n})$ for large $n$, using computations involving only smaller values of $n$. Recall that $f$ is a Gaussian process with mean function $\mu:\mathcal{X}\to\mathbb{R}^{d}$ and covariance kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, denoted $f\sim\mathcal{GP}(\mu,\kappa)$, if for every $x_{1},\dots,x_{n}\in\mathcal{X}$ and $d\geq 1$, $(f(x_{i}))_{i=1}^{n}\sim\mathcal{N}((\mu(x_{i}))_{i=1}^{n},(\kappa(x_{i},x_{j}))_{i,j=1}^{n})$. Our results are founded on the following lemma, which we prove in Appendix G.
Lemma 1.
Let $f:\mathcal{X}\to\mathbb{R}^{d}$ be a zero-mean vector-valued $m$-dimensional Gaussian process with covariance kernel $\kappa$. For each $n\geq 2$, let
$E(n)\coloneqq\mathbb{E}[d^{-\frac{1}{2}}\|f(x_{n})\|^{2}\;|\;f(x_{i})=0\text{ for }i=1,\dots,n-1],$
denote the mean-squared error of fitting the $f$ to the zero function using $x_{1},\dots,x_{n-1}$. Then
$\frac{\mathsf{pdet}(\mathbf{K}_{n})}{\mathsf{pdet}(\mathbf{K}_{n-1})}\leq\left(\frac{d}{r}\right)^{r/2}E(n)^{r},$
where $r$ is the rank of $\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)$. In the case where $\mathrm{rank}(\mathbf{K}_{n})-\mathrm{rank}(\mathbf{K}_{n-1})=d$ this reduces to
$\frac{\mathsf{pdet}(\mathbf{K}_{n})}{\mathsf{pdet}(\mathbf{K}_{n-1})}\leq E(n)^{d},\qquad\mbox{for any }n>1,$
with equality if $d=1$.
Lemma 1 is particularly interesting since it highlights a connection between the determinants of Gram matrices and error curves for GPs. This bounds the effect of adding or removing a datapoint on the determinant in terms of the prior variance of a corresponding GP. Previous theoretical (see Appendix F) and empirical studies (Spigler et al., , 2020; Bahri et al., , 2021; Li et al., , 2023; Barzilai & Shamir, , 2023) have established a power law relationship of the form $E(n)=\Theta(n^{-\xi})$ as $n\to\infty$ for some $\xi>0$. Invoking Lemma 1 in view of this, we propose the following scaling law for determinants, demonstrated in Figure 2.
Figure 2: Demonstration of a scaling law for the ratios of successive determinants $\det(\mathbf{K}_{n})/\det(\mathbf{K}_{n-1})$ (Assumption 1) for the empirical neural tangent kernel Gram matrix of a trained ResNet50 network on CIFAR-10 with $n=$50000$$ datapoints.
Figure 3: Demonstration of the stationarity and second moment behavior (Assumption 2) of the corresponding logarithmic process.
Assumption 1 (Scaling Law).
Assume there exists a constant $C>0$ and exponent $\nu>0$ such that as $n\to\infty$,
$\frac{\det(\mathbf{K}_{n})}{\det(\mathbf{K}_{n-1})}=\frac{C}{n^{\nu}}[1+o_{p}(1)].$ (3)
This assumption allows for accurate estimation of the log-determinants of interest, and an appropriate law of large numbers. However, to construct corresponding confidence intervals, we will need to assume some properties of the error term appearing in Assumption 1. We further impose the mild assumptions of stationarity and bounded variance, demonstrated in Figure 3.
Assumption 2 (Stationarity).
Assume there exists a constants $C>0$ and exponent $\nu>0$ such that the process $\delta_{n}$ satisfying
$\delta_{n-1}\coloneqq\log\left(\frac{\det(\mathbf{K}_{n})}{\det(\mathbf{K}_{n-1})}\right)-\log(Cn^{-\nu}),\quad n=2,3,\dots,$
is stationary, ergodic, has finite second moment $(\mathbb{E}[\delta_{n}^{2}]<+\infty)$, and $\mathbb{E}[\delta_{n}|\delta_{1},\dots,\delta_{n-1}]=0$.
Under Assumption 1 and Assumption 2, we derive an expression for the asymptotic behavior of the normalized log-determinants. The proof is given in Appendix G.
Proposition 1.
For larger numbers of inputs, letting
$L_{n}$ $\coloneqq n^{-1}\mathsf{logdet}(\mathbf{K}_{n})$
$\hat{L}_{n}$ $\coloneqq L_{1}+\left(1-\frac{1}{n}\right)c_{0}-\nu\frac{\log(n!)}{n},$
a law of large numbers (LLN) and central limit theorem (CLT) hold for the log-determinants $L_{n}$:
  • (LLN) Under Assumption 1, there exist constants $c_{0},\nu>0$ such that as $n\to\infty$,
    $L_{n}=\hat{L}_{n}+o_{p}(1).$ (4)
  • (CLT) Under Assumption 2, there exist constants $c_{0},\nu,\sigma>0$ such that as $n\to\infty$,
    $\frac{n}{\sqrt{n-1}}(L_{n}-\hat{L}_{n})\overset{\mathcal{D}}{\to}\mathcal{N}(0,\sigma^{2}).$ (5)
We remark that $\log(n!)$ is $\Theta(n\log(n))$, so the normalized log-determinant $L_{n}$ is $\Theta_{p}(\log(n))$ via (4).
3.2 FLODANCE
As (4) is linear in the unknown parameters $c_{0},\nu$, these parameters can be estimated by linear regression on a precomputed sequence  [^2] This is generally a more stable regression problem than estimating $C$ and $\nu$ directly from (3). $(L_{1},\dots,L_{n})$. This sequence can be computed using a single pass of MEMDET on a subsample of the Gram matrix. After performing linear regression, by extrapolating to larger $n$, the normalized log-determinants of larger NTK Gram matrices can be estimated. The ordering of the data is arbitrary, but will affect the output of the regression task. We refer to this as the Factorial-based Log-Determinant Analysis and Numerical Curve Estimation procedure, or FLODANCE. To improve performance, we allow for a non-asymptotic correction to the exponent $\nu$ as $\nu(n)=\nu_{0}+\sum_{i=1}^{q}\nu_{i}n^{-i}$ (in practice, we find $q=3\sim 6$ to work well). In light of (5) and discussion in Appendix H, for large $n$, we expect that approximately
$y_{n}=c_{0}x_{n,0}+\sum_{i=1}^{q+1}\nu_{i-1}x_{n,i}+\epsilon_{n},\quad\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2}),$
where $y_{n}=\frac{n}{\sqrt{n-1}}(L_{n}-L_{1})$, $x_{n,0}=\sqrt{n-1}$, and $x_{n,i}=\frac{-\log(n!)}{n^{i-1}\sqrt{n-1}}$ for $i=1,...,m+1$ are the covariates. The corresponding numerical procedure is presented in Algorithm 1 and the implementation can be found in LABEL:list:detkit2.
1
Input : Precomputed partial NTK Gram matrix $\mathbf{K}_{n_{s}}$ of size $m_{s}\times m_{s}$ where $m_{s}=n_{s}d$,
Number of model’s output dimension $d$ (i.e., classes or labels)
Total number of datapoints $n$ (i.e., training set size),
Number of datapoint samples $n_{s}$ where $1<n_{s}\leq n$,
Number of burn-in terms $n_{0}$ where $1\leq n_{0}<n_{s}$,
Number of terms in the Laurent series $q$
Output : Estimated normalized log-determinant $\hat{L}_{n}$ of matrix $\mathbf{K}_{n}$ of size $m\times m$ where $m=nd$
2
3
// Denote $\mathbf{K}_{n_{s}[:k,:k]}$ the $k\times k$ principal sub-matrix of $\mathbf{K}_{n_{s}}$
4 Run Algorithm 3 on $\mathbf{K}_{n_{s}}$ to obtain $(\ell_{k})_{k=1}^{m_{s}}$ where $\ell_{k}\leftarrow\mathsf{logabsdet}(\mathbf{K}_{n_{s}[:k,:k]})$
// Normalize log-determinants and record every $d$-th entry in the sequence.
5 Obtain $(L_{j})_{j=n_{0}}^{n_{s}}$ where $L_{j}\leftarrow j^{-1}\ell_{m_{j}}$ and $m_{j}\leftarrow jd$.
6
// Define design matrix $\mathbf{X}\in\mathbf{R}^{(n_{s}-n_{0})\times(q+2)}$ and response vector $\boldsymbol{y}\in\mathbb{R}^{n_{s}-n_{0}}$
7
8for $j=1$ to $n_{s}-n_{0}$ do
9       $n_{j}\leftarrow n_{0}+j$
10       $y_{j}\leftarrow n_{j}(n_{j}-1)^{-\frac{1}{2}}(L_{n_{j}}-L_{n_{0}})$
11      $X_{j,1}\leftarrow(n_{j}-1)^{\frac{1}{2}}$
12      for $i=1$ to $q+1$ do
             $X_{j,i+1}\leftarrow-(n_{j}-1)^{-\frac{1}{2}}n_{j}^{-i+1}\operatorname{log\Gamma}(n_{j}+1)$
              // $\operatorname{log\Gamma}(n+1)=\log(n!)$ is the Log-gamma function.
13
14
15
// Perform linear regression $\boldsymbol{y}=\mathbf{X}\boldsymbol{\beta}+\epsilon$ to estimate coefficients $\boldsymbol{\beta}$
16 $\boldsymbol{\beta}\leftarrow(\mathbf{X}^{\intercal}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{y}$
// Estimate $L_{n}$ at larger value of $n$
17 $(c_{0},\nu_{0},...,\nu_{q})\leftarrow\boldsymbol{\beta}$
18 $\hat{L}_{n}\leftarrow L_{n_{0}}+c_{0}(1-n^{-1})-\sum_{i=1}^{q+1}\nu_{i-1}n^{-i}\operatorname{log\Gamma}(n+1)$
return $\hat{L}_{n}$
Algorithm 1 FLODANCE: Factorial-based Log-Determinant Analysis and Numerical Curve Estimation
In practice, we also observe that a burn-in period may be required to obtain accurate estimates of $c_{0}$ and the $\nu_{i}$ that appear in Proposition 1. Better performance was often achieved in our experiments by discarding the early determinant samples, effectively replacing the $L_{1}$ term appearing in $y_{n}$ with different constants $L_{k+1}$, for a burn-in of length $k$. This appears to be due to the sudden appearance of very small eigenvalues that shift the model fit, and appears to be a consistent phenomena that warrants further investigation. We found that when needed, the burn-in required was always less than 500 terms, verified by cross-validation.
Table 2: Comparison of various approximations of the normalized log-determinant $\hat{L}_{n}$ with the exact computation $L_{n}$ obtained in 64-bit floating-point precision (first row). Values represent averages over five trained networks, with standard deviations in parentheses. Bold values indicate the closest approximation, with the next-best underlined.
Quantity Model Configuration ResNet9 ResNet9 ResNet18 MobileNet Dataset CIFAR-10 CIFAR-10 CIFAR-10 MNIST Subsample Size $n=1000$ $n=2500$ $n=1000$ $n=2500$ $L_{n}$ Direct Computation (64-bit) (Reference) $-0.490$ ($0.20$) $-4.48$ ($0.26$) $-15.5$ ($0.84$) $-81.0$ ($3.15$) Approximation $(\hat{L}_{n})$ Direct Computation (16-bit) $9.00$ ($0.16$) $7.89$ ($0.21$) $-6.32$ ($1.07$) $-26.7$ ($0.85$) Direct Computation (32-bit) $2.32$ ($0.18$) $0.43$ ($0.22$) $-12.1$ ($0.83$) $\boldsymbol{-69.4}$ ($2.00$) Block Diagonal $18.0$ ($0.13$) $18.0$ ($0.17$) $4.77$ ($1.06$) $10.2$ ($1.40$) SLQ $14.2$ ($0.15$) $13.9$ ($0.22$) $-0.66$ ($1.06$) $-12.5$ ($1.37$) Pseudo NTK $-13.9$ ($0.05$) $-16.2$ ($0.15$) $\boldsymbol{-14.89}$ ($0.15$) $-19.9$ ($0.19$) FLODANCE $n_{0}=1,n_{s}=50$ $4.32$ ($0.46$) $1.96$ ($0.69$) $-8.40$ ($1.20$) $-55.0$ ($2.20$) FLODANCE $n_{0}=1,n_{s}=100$ $3.31$ ($0.27$) $0.91$ ($0.54$) $-10.6$ ($0.70$) $-54.0$ ($2.06$) FLODANCE $n_{0}=300,n_{s}=500$ $\boldsymbol{0.53}$ ($0.27$) $\boldsymbol{-3.50}$ ($0.33$) $-13.6$ ($0.43$) $-62.3$ ($1.13$)
4 Numerical Experiments
We are now in a position to test the accuracy of the FLODANCE algorithm for estimation of log-determinants on large-scale problems of interest. The test problems that we consider are NTK matrices corresponding to common deep learning models: ResNet9, ResNet18, and ResNet50 (He et al., , 2016) trained on the CIFAR-10 dataset (Krizhevsky, , 2009), and MobileNet (Howard et al., , 2017) trained on the MNIST dataset (LeCun et al., , 1998). Our experiments are split into two sections: first, the dataset size is reduced in order to enable the matrices to fit into memory on a consumer device, in Section 4.1; and then larger subsamples and full datasets are considered, in Section 4.2.333Experiments in this section were conducted on a desktop-class device with an AMD Ryzen®7 5800X processor, NVIDIA RTX 3080, and 64GB RAM. As a baseline, we employ MEMDET to compute the relevant quantities, where the accuracy is limited only by numerical precision.
4.1 Smaller Data Sets
In order to demonstrate the scaling laws for ratios of successive determinants of NTK Gram matrices, a ResNet50 model was trained on a subset of 1000 datapoints from the CIFAR-10 dataset. The resulting matrix $\mathbf{K}_{1000}$ is of size $10000$\times$10000$$. In Figure 2, we plot exact ratios $\det(\mathbf{K}_{n})/\det(\mathbf{K}_{n-1})$, with the blue line representing the line of best fit under the scaling law Assumption 1.
For baseline comparison to existing techniques, we compared FLODANCE to approximations of both the matrices themselves and their log-determinants. For the matrix approximations, we consider a block-diagonal approximation ignoring between-data correlations, as well as the pseudo-NTK matrix studied in (Mohamadi et al., , 2023). As an approximate log-determinant technique, we consider the stochastic Lanczos quadrature, often regarded as the current state-of-the-art in log-determinant estimation of large matrices (Gardner et al., , 2018). We further compare the exact methods across 16-, 32- and 64-bit (treated as exact) floating-point precision, in order to compare the accuracy of our extrapolation method against memory-saving mixed-precision calculations.
To this end, ResNet9 was trained on a subsets of 1000 and 2500 images from CIFAR-10, and ResNet18 on 1000 datapoints. MobileNet was also trained on a 2500 image subset of the MNIST dataset. A comparison of the different methods is presented in Table 2. We see that all existing NTK and log-determinant approximation techniques perform poorly when compared with 16- and 32-bit mixed-precision calculations. On the other hand, the FLODANCE estimates that contained a burn-in phase consistently either outperformed the mixed-precision approximations or were competitive. When no burn-in was used, FLODANCE still outperformed the approximation methods. This suggests that at this scale, the error in the scaling law approximation is less than the mixed-precision errors discussed in Section 2.1. Extrapolating determinants based on expected behavior can bypass numerical issues at scale. Comparisons for ResNet9 trained on 2500 images from the CIFAR-10 dataset are visualized in Figure 4. FLODANCE estimates consistently outperform the competing methods.
Figure 4: Comparison of log-determinant accuracy for NTKs of ResNet9 trained on CIFAR-10, measured by absolute error, across a variety of approximation techniques for matrices of different sizes. Means across five trained networks are displayed, with shaded regions depicting one standard deviation.
Figure 5: Log-determinant $\ell_{n}$ for $n=1,\dots,$50000$$, corresponding to $m\times m$ NTK submatrices where $m=nd$ and $d=10$, from ResNet9 (a) and ResNet50 (b) trained on CIFAR-10 with $50000$ datapoints. Values are computed using MEMDET (Algorithm 3) with LDL decomposition (black curves, overlaid by colored curves). FLODANCE (Algorithm 1) is fitted in a small region (shaded gray) and extrapolated over a much larger interval. The yellow curve in the interval $(n_{0},n_{s})=(10^{2},5\times 10^{3})$ represents the fit, while the red curve in $(n_{s},n)=(5\times 10^{3},5\times 10^{4})$ shows the extrapolation. The blue curves, corresponding to the right axis in each panel shows the relative error, reaching impressive 0.05% in (a) and 0.02% in (b).
4.2 Larger Data Sets
In our final experiment, we evaluate NTK matrices at an unprecedented scale, where exact determinant computation has never been previously reported. Due to memory constraints, these matrices cannot be stored explicitly, making MEMDET essential for obtaining ground truth values.
We consider two large-scale NTK matrices: $\mathbf{K}_{$50000$}$, a dense matrix of size $500000$\times$500000$$, for ResNet50 trained on CIFAR-10, and an identical-sized NTK for ResNet9 trained on CIFAR-10. At this scale, double-precision arithmetic is infeasible due to memory and computational constraints, making 32-bit computation the only viable direct baseline. However, even in 32-bit, these matrices require nearly a terabyte of storage.
As shown in Figure 5, FLODANCE with $q=6$, $n_{0}=100$, and $n_{s}=5000$ achieves an absolute error of just $0.05\%$ for $L_{$50000$}$ on ResNet9, reducing computation time by a factor of $(n/n_{s})^{3}=1000$. Similarly, for ResNet50, FLODANCE with $q=4$, $n_{0}=100$, and $n_{s}=5000$ achieves an absolute error of $0.02\%$ with the same speedup. In contrast, SLQ exhibited a relative error of 55% (see Table 1, also LABEL:list:imate for implementation). Given their poor performance on smaller datasets, pseudo-NTK and block-diagonal approximations are omitted.
This experiment represents the first exact computation of an NTK determinant at this scale, establishing a new benchmark for large-scale log-determinant estimation.
5 Conclusion
The calculation of large matrix log-determinants is a commonly encountered but often avoided problem when considering statistical and machine learning problems at scale. A number of techniques have previously been proposed to circumvent explicit computation, typically relying upon stochastic approximations. However, in many problems of interest the sheer size of the matrices, combined with their highly ill-conditioned nature, make not only approximation a difficult task, but forming the matrix itself to provide a baseline becomes computationally intractable. We have addressed this problem on two fronts. On the one hand, we defined MEMDET, a hierarchical memory-constrained algorithm for log-determinant computation, with different versions for general, symmetric, and symmetric positive-definite matrices. On the other hand, we derived neural scaling laws for large kernel matrices, and introduced FLODANCE, a procedure for accurate extrapolation of log-determinants from small subsets of the data. The high level of speed and accuracy of our methods opens the door for routine computation of the interpolating information criteria and related diagnostic tools to enable principled model selection within deep learning frameworks (Hodgkinson et al., 2023b, ).
The ability to accurately compute and estimate matrices of this size further provides fascinating insights into the behavior of the NTKs that we considered in our experiments, which considered square matrices of the size up to $500000$. Further, the memory constrained algorithms we considered can be applied to other classes of matrices (Nguyen & Vu, , 2014; Cai et al., , 2015), where they can be expected to unlock similar insights into their scaling behavior. In terms of further computational tools, we hope for the development of blockwise decompositions of large scale Jacobian matrices, in order to circumvent the need to explicitly calculate $\mathbf{J}\mathbf{J}^{\intercal}$.
Acknowledgments. LH is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144). FR was partially supported by the Australian Research Council through an Industrial Transformation Training Centre for Information Resilience (IC200100022). MWM would like to acknowledge the DOE, IARPA, NSF, and ONR for providing partial support of this work.
Appendix A Related Works
The study of block decomposition methods in numerical linear algebra has a long history. Classical texts such as Golub & Van Loan, (2013) and Dongarra et al., 1998b provide foundational discussions on block LU, block Cholesky, and LDL decompositions, detailing their computational advantages and numerical properties. These methods have been widely used to improve computational efficiency, particularly in high-performance computing (HPC) settings, where recursive block LU (Golub & Van Loan, , 2013, Section 3.2.11), parallel LU (Golub & Van Loan, , 2013, Section 3.6), and block Cholesky (Golub & Van Loan, , 2013, Section 4.2.9) play a central role in large-scale matrix computations.
Beyond theoretical foundations, numerous works have focused on efficient implementations of block factorizations, particularly for parallel architectures. Stark & Beris, (1992) optimized block LU decomposition for hierarchical distributed memory, aiming to improve data locality while maintaining parallel efficiency. Dongarra et al., (1979) pioneered high-performance implementations of block factorizations, laying the groundwork for modern HPC systems. More recent studies, such as Galoppo et al., (2005) and Barrachina et al., (2008), have extended block LU methods to GPU-based environments, leveraging parallelism but still assuming that intermediate submatrices fit in memory. While these approaches optimize performance in parallel settings, they do not address the challenge of computing factorizations when the full matrix size far exceeds available RAM. Traditional block methods typically assume that at least some intermediate submatrices can reside in memory, whereas our method (MEMDET) explicitly operates under constrained memory settings, utilizing an out-of-core hierarchical block processing approach.
To address the issue of matrices exceeding main memory capacity, Dongarra et al., 1998a introduced concepts for parallel out-of-core LU factorization, focusing on efficient data movement between disk and memory. While their work demonstrates how out-of-core computations can be applied to LU factorization, their approach does not extend to log-determinant computations or hierarchical block-wise processing. Similarly, studies on many-core architectures (Venetis & Gao, , 2009) and hierarchical memory-aware LU factorizations (Demmel et al., , 1993) have improved computational efficiency, but none have been designed specifically for computing log-determinants under extreme memory constraints, making our approach distinct.
Appendix B Memory and Computation Challenges of NTK Matrices
The empirical NTK serves as a motivating example throughout this work, as it encapsulates many of the computational challenges associated with large-scale matrix operations. Several software packages have been developed to compute NTK Gram matrices for different neural architectures using automatic differentiation frameworks (Novak et al., , 2022; Engel et al., , 2022). However, the full formation of these matrices remains computationally prohibitive, even on commonly used benchmark datasets.
Table B.1 presents the storage requirements for NTK matrices corresponding to various datasets, highlighting their enormous size. For instance, even CIFAR-10 requires terabytes of storage, while ImageNet-1k exceeds exabytes, making full NTK computation infeasible for most practical applications. Despite its theoretical importance, the NTK Gram matrix is rarely used as a practical tool, with approximations often employed to mitigate computational and memory constraints. Minibatching is one common strategy, and batch-wise NTK approximations have been explored for model selection (Immer et al., , 2023), yet extending these estimates to full datasets remains an open challenge. Alternative approximation techniques (Mohamadi et al., , 2023) have been proposed, but their convergence is only guaranteed in spectral norm, limiting their ability to capture the full spectrum of the NTK. In contrast, the log-determinant—a key quantity in this work—encodes information from the entire eigenvalue distribution, making its computation particularly demanding.
Table B.1: Memory requirements (for different data type precisions) to store empirical NTK matrices of common datasets.
Matrix Size
Dataset Training Set Classes float16 float32 float64
CIFAR-10 $50000$ 10 0.5 TB 1.0 TB 2.0 TB
MNIST $60000$ 10 0.72 TB 1.5 TB 2.9 TB
SVHN $73257$ 10 1.1 TB 2.2 TB 4.2 TB
ImageNet-1k $1281167$ 1000 $3282778$ TB $6565556$ TB $13131111$ TB
Beyond storage limitations, the computation time for NTK matrices also presents a major challenge. Table B.2 provides estimated compute times for NTK formation across various models and floating-point precisions on an NVIDIA H100 GPU. Even for relatively small datasets like CIFAR-10, NTK computation is expensive, with higher-precision calculations significantly increasing runtime. For large architectures such as ResNet101, double-precision NTK computation can require over a thousand hours, making exact evaluations impractical without algorithmic improvements like those introduced in this work.
Table B.2: Estimated compute time (in hours using NVIDIA H100 GPU) for NTK matrix computation.
Compute Time (hrs)
Dataset Model float16 float32 float64
MNIST MobileNet 6 25 50
CIFAR-10 ResNet9 6 24 70
ResNet18 14 63 65
ResNet50 37 177 297
ResNet101 107 442 1178
Appendix C Implementation of MEMDET Algorithm
The pseudo-code of MEMDET algorithm is given in Algorithm 2 (for generic matrices), Algorithm 3 (for symmetric matrices), and Algorithm 4 (for symmetric positive-definite matrices), which computes the log-determinant of the matrix $\mathbf{M}$. We note that the log-determinant of all sub-matrices $\mathbf{M}_{[:k,:k]}$, or their row/column permutations for $k=1,\dots,m$, can also be readily computed. For instance, for symmetric and positive-definite matrices using Cholesky decomposition $\mathbf{M}=\mathbf{L}\mathbf{L}^{\intercal}$, with lower-triangular matrix $\mathbf{L}$, the logarithm of its positive determinant is $\mathsf{logdet}(\mathbf{M}_{[:k,:k]})=2\sum_{i=1}^{k}\log(L_{ii})$.
The memory requirements of MEMDET are determined by a user-defined parameter, allowing it to run on any system regardless of available memory. For an $m\times m$ matrix, the algorithm partitions the data into an $n_{b}\times n_{b}$ grid of blocks, each of size $b\times b$, where $b=1+\lfloor(m-1)/n_{b}\rfloor$. The computation requires either 3 or 4 concurrent blocks in memory: for $n_{b}=2$, only 3 blocks are needed, requiring $3b^{2}\beta$ bytes, while for $n_{b}>2$, 4 blocks are required, increasing the memory usage to $4b^{2}\beta$ bytes, where $\beta$ is the number of bytes per floating point.
Given a maximum memory limit $c$ (in bytes), the optimal number of blocks $n_{b}$ is determined by the parameter $r=m\sqrt{\beta/c}$, with the following selection criteria:
  • If $r\leq 1$, the entire matrix fits in memory, so $n_{b}=1$.
  • If $r\leq\frac{2}{\sqrt{3}}$, three blocks fit in memory, so $n_{b}=2$.
  • Otherwise, $n_{b}=\lceil 2r\rceil$.
1
2
3
Input : Matrix $\mathbf{M}$ of size $m\times m$, // stored on disk, may not be loaded on memory
Maximum memory $c$ in bytes
Output : $\ell$: logarithm of the absolute value of the determinant ($\mathsf{logabsdet})$ of $\mathbf{M}$,
$\sigma$: sign of the determinant of $\mathbf{M}$
4
$r\leftarrow m\sqrt{\beta/c}$
  // $\beta$: number of bytes per floating-point
5
6 if $r\leq 1$ then  $n_{b}\leftarrow 1$ // $n_{b}$: number of row/column blocks, making $n_{b}\times n_{b}$ grid of blocks.
7 else if $r\leq 2/\sqrt{3}$ then  $n_{b}\leftarrow 2$
8 else  $n_{b}\leftarrow\lceil 2r\rceil$
9
$b\leftarrow 1+\lfloor(m-1)/n_{b}\rfloor$
  // Size of each block is at most $b\times b$
$\ell\leftarrow 0$
  // Accumulates log-abs-determinant of diagonal blocks
$\sigma\leftarrow 1$
  // Keeps track of the parity of matrix
10
11
// Allocate memory for block matrices
12 Allocate memory for $b\times b$ matrix $\mathbf{A}$
13
14if $n_{b}>1$ then  Allocate memory for $b\times b$ matrices $\mathbf{B},\mathbf{C}$
15
16if $n_{b}>2$ then  Allocate memory for $b\times b$ matrix $\mathbf{S}$
17
// Create scratchpad space on disk, large enough to store $n_{b}(n_{b}-1)-1$ blocks
18 if $n_{b}>2$ then  Allocate empty file of the size $(m(m-b)-b^{2})\beta$ bytes
19
// Recursive iterations over diagonal blocks
20 for $k=1$ to $n_{b}$ do
21       if $k=1$ then  $\mathbf{A}\leftarrow\mathbf{M}_{kk}$ // Load from input array on disk
       $\mathbf{A}\leftarrow\mathbf{P}\mathbf{L}\mathbf{U}$
        // In-place LU decomposition with pivoting (written to $\mathbf{A}$)
22       $\ell\leftarrow\ell+\mathsf{logabsdet}(\mathbf{U})$
23      $\sigma\leftarrow\sigma\;\mathsf{sgn}(\mathbf{P})\;\mathsf{sgn}(\mathbf{U})$
24       if $k<n_{b}$ then
             // Iterate over row of blocks from bottom upward
25             for $i=n_{b}$ to $k+1$ step $-1$ do
                   $\mathbf{C}\leftarrow\mathbf{M}_{ik}^{\intercal}$
                    // Load from disk (from input array if $k=1$ or from scratchpad if $k>1$)
                   $\mathbf{C}\leftarrow\mathbf{U}^{-\intercal}\mathbf{C}$
                    // Solve upper triangular system in-place
26
                   // Iterate over column of blocks in alternating directions per row
27                   if $i-k$ is even then  $(j_{\operatorname{start}},j_{\operatorname{end}})\leftarrow(k+1,n_{b})$
28                   else  $(j_{\operatorname{start}},j_{\operatorname{end}})\leftarrow(n_{b},k+1)$
29
30                   for $j=j_{\operatorname{start}}$ to $j_{\operatorname{end}}$ step $(-1)^{i-k}$ do
                         // Load $\mathbf{B}$ from disk (input array if $k=1$ and $i=n_{b}$, otherwise from scratchpad)
31                         if $i=n_{b}$ or $j\neq j_{\operatorname{start}}$ then  $\mathbf{B}\leftarrow\mathbf{M}_{kj}$
32
33                         if $i=n_{b}$ then
34                               $\mathbf{B}\leftarrow\mathbf{L}^{-1}\mathbf{P}^{\intercal}\mathbf{B}$ // Solve lower triangular system in-place
35                              if $n_{b}-k>2$ or $j\neq j_{\operatorname{end}}$ then  $\mathbf{M}_{kj}\leftarrow\mathbf{B}$ // Write to disk on scratchpad
36
37
38                         if $i=k+1$ and $j=k+1$ then
                               $\mathbf{A}\leftarrow\mathbf{M}_{ij}$
                                // Load from disk (input array if $k=1$ or scratchpad if $k>1$)
                               $\mathbf{A}\leftarrow\mathbf{A}-\mathbf{C}^{\intercal}\mathbf{B}$
                                // Compute Schur complement
39
40                        else
                               $\mathbf{S}\leftarrow\mathbf{M}_{ij}$
                                // Load from disk (input array if $k=1$ or scratchpad if $k>1$)
                               $\mathbf{S}\leftarrow\mathbf{S}-\mathbf{C}^{\intercal}\mathbf{B}$
                                // Compute Schur complement
                               $\mathbf{M}_{ij}\leftarrow\mathbf{S}$
                                // Write to disk on scratchpad
41
42
43
44
45
46
47 return $\ell$, $\sigma$
Algorithm 2 MEMDET: Constrained-Memory Comp. of Log-Det (Case I: Generic Matrix)
1
2
1
2
Input : Symmetric matrix $\mathbf{M}$ of size $m\times m$, // stored on disk, may not be loaded on memory
Maximum memory $c$ in bytes
Output : $\boldsymbol{\pi}\coloneqq(\pi_{q})_{q=1}^{m}$: a permutation of $\{1,\dots,m\}$ // permutations induced by LDL decomposition
$\boldsymbol{\ell}\coloneqq(\ell_{q})_{q=1}^{m}$: $\ell_{q}\coloneqq\mathsf{logabsdet}(\mathbf{M}_{[\mathcal{I}_{q},\mathcal{I}_{q}]})$ with the index set $\mathcal{I}_{q}\coloneqq(\pi_{1},\dots,\pi_{q})$ // $\log|\det(\cdot)|$
$\boldsymbol{\sigma}\coloneqq(\sigma_{q})_{q=1}^{m}$: $\sigma_{q}\coloneqq\mathsf{sgn}(\det(\mathbf{M}_{[\mathcal{I}_{q},\mathcal{I}_{q}]}))$ // Sign of $\det(\cdot)$
3
$r\leftarrow m\sqrt{\beta/c}$
  // $\beta$: number of bytes per floating-point
4
5 if $r\leq 1$ then  $n_{b}\leftarrow 1$ // $n_{b}$: number of row/column blocks, making $n_{b}\times n_{b}$ grid of blocks.
6 else if $r\leq 2/\sqrt{3}$ then  $n_{b}\leftarrow 2$ else  $n_{b}\leftarrow\lceil 2r\rceil$
7
$b\leftarrow 1+\lfloor(m-1)/n_{b}\rfloor$
  // Size of each block is at most $b\times b$
Initialize arrays $\boldsymbol{d}\in\mathbb{R}^{m}$ and $\boldsymbol{\pi}\in\{1,\dots,m\}^{m}$
  // Hold diagonals and permutations, respectively
8
9
// Allocate memory for block matrices
10 Allocate memory for $b\times b$ matrix $\mathbf{A}$
11
12if $n_{b}>1$ then  Allocate memory for $b\times b$ matrices $\mathbf{B},\mathbf{C}$
13
14if $n_{b}>2$ then  Allocate memory for $b\times b$ matrix $\mathbf{S}$
15
16if $n_{b}>1$ then  Define pointers $\mathbf{B}_{\star},\mathbf{C}_{\star}$ // Used for swapping memory; $(\mathbf{B}_{\star},\mathbf{C}_{\star})$ will refer to $(\mathbf{B},\mathbf{C})$ or $(\mathbf{C},\mathbf{B})$
17
// Create scratchpad space on disk, large enough to store $n_{b}(n_{b}+1)/2-4$ blocks
18 if $n_{b}>2$ then  Allocate empty file of the size $(m(m+b)/2-4b^{2})\beta$ bytes
19
// Recursive iterations over diagonal blocks
20 for $k=1$ to $n_{b}$ do
21       if $k=1$ then  $\mathbf{A}\leftarrow\mathbf{M}_{kk}$ // Load from input array on disk
22
      $\mathbf{A}\leftarrow\mathbf{P}\mathbf{L}\mathbf{D}\mathbf{L}^{\intercal}\mathbf{P}^{\intercal}$
        // In-place LDLT decomposition with pivoting (written to $\mathbf{A}$)
23
24      $\boldsymbol{d}_{[1+(k-1)b:kb]}\leftarrow\mathsf{diag}(\mathbf{D})$ // Accumulate diagonals of $\mathbf{D}$ to $\boldsymbol{d}$
25      $\boldsymbol{\pi}_{[1+(k-1)b:kb]}\leftarrow(k-1)b+\mathsf{permutation(\mathbf{P})}$ // Accumulate permutation indices
26       if $k<n_{b}$ then
             // Iterate over column of blocks backward (right to left)
27             for $j=n_{b}$ to $k+1$ step $-1$ do
28                   if $n_{b}-j$ is even then  $(\mathbf{B}_{\star},\mathbf{C}_{\star})=(\mathbf{B},\mathbf{C})$ else  $(\mathbf{B}_{\star},\mathbf{C}_{\star})=(\mathbf{C},\mathbf{B})$ // swap $\mathbf{B}$ and $\mathbf{C}$ memories
29
30                   if $j=n_{b}$ then
                         $\mathbf{B}_{\star}\leftarrow\mathbf{M}_{kj}$
                          // Load from disk (input array if $k=1$ or from scratchpad if $k>1$)
31
32                        $\mathbf{B}_{\star}\leftarrow\mathbf{L}^{-1}\mathbf{P}^{\intercal}\mathbf{B}_{\star}$ // Solve lower-triangular system in-place
33
34                   $\mathbf{C}\leftarrow\mathbf{B}_{\star}$ // Deep copy of the memory pointed by $\mathbf{B}_{\star}$ to memory pointed by $\mathbf{C}_{\star}$
35                  $\mathbf{B}_{\star}\leftarrow\mathbf{D}^{-1}\mathbf{B}_{\star}$
                   // Processing order of rows: first process row $j$, then from row $k+1$ downward to $j-1$
36                   $\mathcal{R}\leftarrow(j,k+1,k+2,\dots,j-2,j-1)$
37                   for $i=\mathcal{R}(1)$ to $\mathcal{R}(j-k)$ do
38                         if $i\neq j$ then  $\mathbf{C}_{\star}\leftarrow\mathbf{M}_{ki}$ // Load disk (input array if $k=1$, $j=n_{b}$, otherwise scratchpad)
39
40                         if $j=n_{b}$ then
41                               $\mathbf{C}_{\star}\leftarrow\mathbf{L}^{-1}\mathbf{P}^{\intercal}\mathbf{C}_{\star}$ // Solve lower triangular system in-place
42                              if $n_{b}>2$ and $i<j-1$ then  $\mathbf{M}_{ki}\leftarrow\mathbf{C}_{\star}$ // Write to disk on scratchpad
43
44
45                         if $i=k+1$ and $j=k+1$ then
                               $\mathbf{A}\leftarrow\mathbf{M}_{ij}$
                                // Load from disk (input array if $k=1$ or scratchpad if $k>1$)
                               $\mathbf{A}\leftarrow\mathbf{A}-\mathbf{C}_{\star}^{\intercal}\mathbf{B}_{\star}$
                                // Compute Schur complement
46
47                        else
                               $\mathbf{S}\leftarrow\mathbf{M}_{ij}$
                                // Load from disk (input array if $k=1$ or scratchpad if $k>1$)
                               $\mathbf{S}\leftarrow\mathbf{S}-\mathbf{C}_{\star}^{\intercal}\mathbf{B}_{\star}$
                                // Compute Schur complement
                               $\mathbf{M}_{ij}\leftarrow\mathbf{S}$
                                // Write to disk on scratchpad
48
49
50
51
52
53
54
55$(\ell_{0},\sigma_{0})\leftarrow(0,1)$
56for $q=1$ to $m$ do  $(\ell_{q},\sigma_{q})\leftarrow(\ell_{q-1}+\log(|d_{q}|),\sigma_{q-1}\,\mathsf{sgn}(d_{q}))$ // $d_{q}$ is the $q$-th element of $\boldsymbol{d}$
57
58 return $\boldsymbol{\pi},\boldsymbol{\ell}$, $\boldsymbol{\sigma}$
Algorithm 3 MEMDET: Constrained-Memory Comp. of Log-Det (Case II: Symmetric Matrix)
1
2
1
2
Input : Symmetric positive-definite matrix $\mathbf{M}$ of size $m\times m$, // stored on disk, not on memory
Maximum memory $c$ in bytes
Output : $\boldsymbol{\ell}\coloneqq(\ell_{q})_{q=1}^{m}$: $\ell_{q}\coloneqq\mathsf{logdet}(\mathbf{M}_{[:q,:q]})$
3
$r\leftarrow m\sqrt{\beta/c}$
  // $\beta$: number of bytes per floating-point
4
5 if $r\leq 1$ then  $n_{b}\leftarrow 1$ // $n_{b}$: number of row/column blocks, making $n_{b}\times n_{b}$ grid of blocks.
6 else if $r\leq 2/\sqrt{3}$ then  $n_{b}\leftarrow 2$ else  $n_{b}\leftarrow\lceil 2r\rceil$
7
$b\leftarrow 1+\lfloor(m-1)/n_{b}\rfloor$
  // Size of each block is at most $b\times b$
Initialize array $\boldsymbol{d}\in\mathbb{R}^{m}$
  // Holds diagonals
8
9
// Allocate memory for block matrices
10 Allocate memory for $b\times b$ matrix $\mathbf{A}$
11
12if $n_{b}>1$ then  Allocate memory for $b\times b$ matrices $\mathbf{B},\mathbf{C}$
13
14if $n_{b}>2$ then  Allocate memory for $b\times b$ matrix $\mathbf{S}$
15
16if $n_{b}>1$ then  Define pointers $\mathbf{B}_{\star},\mathbf{C}_{\star}$ // Used for swapping memory; $(\mathbf{B}_{\star},\mathbf{C}_{\star})$ will refer to $(\mathbf{B},\mathbf{C})$ or $(\mathbf{C},\mathbf{B})$
17
// Create scratchpad space on disk, large enough to store $n_{b}(n_{b}+1)/2-4$ blocks
18 if $n_{b}>2$ then  Allocate empty file of the size $(m(m+b)/2-4b^{2})\beta$ bytes
19
// Recursive iterations over diagonal blocks
20 for $k=1$ to $n_{b}$ do
21       if $k=1$ then  $\mathbf{A}\leftarrow\mathbf{M}_{kk}$ // Load from input array on disk
22
      $\mathbf{A}\leftarrow\mathbf{L}\mathbf{L}^{\intercal}$
        // In-place Cholesky decomposition (written to $\mathbf{A}$)
23
24      $\boldsymbol{d}_{[1+(k-1)b:kb]}\leftarrow\mathsf{diag}(\mathbf{L})$ // Accumulate diagonals of $\mathbf{L}$ to $\boldsymbol{d}$
25       if $k<n_{b}$ then
             // Iterate over column of blocks backward (right to left)
26             for $j=n_{b}$ to $k+1$ step $-1$ do
27                   if $n_{b}-j$ is even then  $\mathbf{B}_{\star}=\mathbf{B}$ else  $\mathbf{B}_{\star}=\mathbf{C}$ // Alternate pointer $\mathbf{B}_{\star}$ to switch between $\mathbf{B}$ and $\mathbf{C}$
28
29                   if $j=n_{b}$ then
                         $\mathbf{B}_{\star}\leftarrow\mathbf{M}_{kj}$
                          // Load from disk (input array if $k=1$ or from scratchpad if $k>1$)
30                         $\mathbf{B}_{\star}\leftarrow\mathbf{L}^{-1}\mathbf{B}_{\star}$ // Solve lower-triangular system in-place
31
                   // Processing order of rows: first process row $j$, then from row $k+1$ downward to $j-1$
32                   $\mathcal{R}\leftarrow(j,k+1,k+2,\dots,j-2,j-1)$
33                   for $i=\mathcal{R}(1)$ to $\mathcal{R}(j-k)$ do
34                         if i = j then  $\mathbf{C}_{\star}=\mathbf{B}_{\star}$ // Shallow copy of pointer $\mathbf{C}_{\star}$ pointing to $\mathbf{B}_{\star}$
35                         else
36                               if $n_{b}-j$ is even then  $\mathbf{C}_{\star}=\mathbf{C}$ else  $\mathbf{C}_{\star}=\mathbf{B}$ // Alternate pointer $\mathbf{C}_{\star}$ between $\mathbf{C}$ and $\mathbf{B}$
37
38                              $\mathbf{C}_{\star}\leftarrow\mathbf{M}_{ki}$ // Load $\mathbf{C}_{\star}$ from disk (input array if $k=1$, $j=n_{b}$, otherwise scratchpad)
39                               if $j=n_{b}$ then
40                                     $\mathbf{C}_{\star}\leftarrow\mathbf{L}^{-1}\mathbf{C}_{\star}$ // Solve lower triangular system in-place
41                                    if $n_{b}>2$ and $i<j-1$ then  $\mathbf{M}_{ki}\leftarrow\mathbf{C}_{\star}$ // Write to disk on scratchpad
42
43
44
45                         if $i=k+1$ and $j=k+1$ then
                               $\mathbf{A}\leftarrow\mathbf{M}_{ij}$
                                // Load from disk (input array if $k=1$ or scratchpad if $k>1$)
                               $\mathbf{A}\leftarrow\mathbf{A}-\mathbf{C}_{\star}^{\intercal}\mathbf{B}_{\star}$
                                // Compute Schur complement
46
47                        else
                               $\mathbf{S}\leftarrow\mathbf{M}_{ij}$
                                // Load from disk (input array if $k=1$ or scratchpad if $k>1$)
                               $\mathbf{S}\leftarrow\mathbf{S}-\mathbf{C}_{\star}^{\intercal}\mathbf{B}_{\star}$
                                // Compute Schur complement
                               $\mathbf{M}_{ij}\leftarrow\mathbf{S}$
                                // Write to disk on scratchpad
48
49
50
51
52
53
54
[…]
56for $q=1$ to $m$ do  $\ell_{q}\leftarrow\ell_{q-1}+2\log(d_{q})$ // $d_{q}$ is the $q$-th element of $\boldsymbol{d}$
57
58 return $\boldsymbol{\ell}$
Algorithm 4 MEMDET: Constrained-Memory Comp. of Log-Det (Case III: Symmetric Positive-Definite Matrix)
Appendix D Optimal Sequence of Processing of Blocks
It is important to select an ordering of the blocks to minimize data transfer between disk and memory. The order in which the block $\mathbf{M}_{ij}$ is processed should be chosen to minimize the reading of the blocks $\mathbf{B}$ (corresponding to the index $j$) and $\mathbf{C}$ (corresponding to the index $i$). Ideally, from processing one block to the next, one should update only one of the matrices $\mathbf{B}$ or $\mathbf{C}$, but not both, to reuse one of the blocks already loaded in memory. We formulate this problem of finding the optimal sequence of blocks as follows for the case of LDL/Cholesky decomposition at the $k$-th stage of the algorithm. The case for LU decomposition can be formulated similarly.
Let $G(V,E)$ denote a complete undirected graph with vertices $V\coloneqq\{k+1,\dots,n_{b}\}$, where $E$ is the set of all possible edges $e=(v,u)$ between the vertices $u,v\in V$. Each vertex in $V$ represents the event of loading one of the blocks $\mathbf{B}\leftarrow\mathbf{M}_{kj}$ or $\mathbf{C}\leftarrow\mathbf{M}_{ik}$, $i,j=k+1,\dots,n_{b}$. Each edge in $E$ corresponds to the event of processing the block $\mathbf{M}_{ij}$. At the $k$-th stage of the algorithm, eventually, all blocks $i,j=k+1,\dots,n_{b}$ will be processed, so $E$ consists of all edges of a complete graph, including self-loops, with $|E|=\frac{|V|(|V|+1)}{2}$. To illustrate this concept, consider the example in Figure D.1. The left panel depicts the $k$-th iteration of the algorithm for a symmetric matrix, where the matrices $\mathbf{B}$ and $\mathbf{C}$ have four blocks to choose from the set $V=\{v_{1},v_{2},v_{3},v_{4}\}$. The corresponding graph $G$ is shown in the middle panel of the figure.
[…]
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$e_{2}$$e_{3}$$e_{4}$$e_{6}$$e_{7}$$e_{9}$$e_{1}$$e_{5}$$e_{8}$$e_{10}$
Figure D.1: Left: Example of the processing order of blocks for a symmetric matrix at the $k$-th hierarchical step. In this step, to process $\mathbf{S}\leftarrow\mathbf{M}_{ij}$, the memory blocks $\mathbf{B}$ and $\mathbf{C}$ are selected from the set $V=\{v_{1},v_{2},v_{3},v_{4}\}$. Middle: The corresponding complete graph $G(V,E)$. Right: The corresponding line graph $L(G)$, with one possible Hamiltonian path highlighted in red, starting from the node $e_{10}$ and ending at the node $e_{1}$.
The goal is to select an ordered sequence $(e_{p})$, $p=1,\dots,|E|$ of edges such that each two consecutive edges $e_{p}$ and $e_{p+1}$ in the sequence share a common vertex. This ensures that from processing one block to the next, only one of $\mathbf{B}$ or $\mathbf{C}$ is updated, while at least one block is reused from the previous step.
To find such a sequence of edges, we define $L(G)$, the line graph of $G$ (also known as the edge-to-vertex dual), where each vertex of $L(G)$ represents an edge of $G$. Two vertices in $L(G)$ are connected if and only if their corresponding edges in $G$ share a common vertex. Consequently, any Hamiltonian path in $L(G)$ defines an ordered sequence of edges that fulfills our requirement.
As illustrated in Figure D.1, the right panel depicts the line graph of the given graph shown in the middle panel, with a possible Hamiltonian path highlighted in red. This path directly translates to the processing order of blocks shown in the left panel. Notably, all valid Hamiltonian paths must terminate at the node $e_{1}$, representing the block $\mathbf{M}_{k+1,k+1}$. This specific end point is crucial as it allows for a seamless transition to the next iteration (i.e., the $k+1$ iteration) without the need to explicitly load the matrix $\mathbf{A}$, as it would already be available in memory from the last processing block of the $k$-th iteration when $\mathbf{S}\leftarrow\mathbf{M}_{k+1,k+1}$ was processed.
Given that $G$ is complete and therefore Hamiltonian, it follows that its line graph $L(G)$ is also Hamiltonian. This implies the existence of at least one (but possibly many) Hamiltonian paths. Crucially, all Hamiltonian paths in $L(G)$ have the same length. Consequently, any sequence of blocks derived from a Hamiltonian path constitutes an optimal solution to our problem. Thus, the block sequence presented in Figure 1 is equivalent in optimality to any other sequence obtainable from a Hamiltonian path.
The same problem can be formulated for the block LU decomposition, with the modification that $G$ is a complete and balanced bipartite graph $G(V,V,E)$; however, the same logic and conclusion follow.
Appendix E Complexity and Performance Analysis of MEMDET
E.1 Computational Complexity
Table E.1 provides a detailed breakdown of the computational complexity of MEMDET for generic matrices (second column, using LU decomposition) and symmetric matrices (third column, using LDL or Cholesky decomposition). The operations are categorized into matrix decomposition, solving triangular systems, and matrix multiplications used to form Schur complements. Each operation’s complexity is given in terms of the number of times it is performed and the FLOP count per operation. In this analysis, a single floating-point fused multiply-add (FMA) operation is counted as one FLOP. The table reports a single complexity column for symmetric matrices, covering both LDL and Cholesky decompositions; while LDL involves additional operations such as row permutations and handling the diagonal matrix $\mathbf{D}$, these steps are excluded from the FLOP count, as their computational cost is negligible compared to the dominant terms.
Table E.1: Breakdown of computational complexity for MEMDET. The table presents the number of operations performed and the FLOP count per operation for generic matrices (LU decomposition) and symmetric matrices (LDL or Cholesky decomposition). The last row shows the total complexity, which remains independent of the number of blocks $n_{b}$.
Generic Matrix Symmetric (Positive-Definite) Matrix Operation Num. Operations FLOPs per Operation Num. Operations FLOPs per Operation Matrix Decomposition $n_{b}$ $\frac{1}{3}b^{3}-\frac{1}{2}b^{2}+\frac{1}{6}b$ $n_{b}$ $\frac{1}{6}b^{3}-\frac{1}{4}b^{2}+\frac{1}{12}b$ Solve Lower Triangular System $\frac{1}{2}n_{b}^{2}-\frac{1}{2}n_{b}$ $\frac{1}{2}b^{3}-\frac{1}{2}b^{2}$ $\frac{1}{2}n_{b}^{2}-\frac{1}{2}n_{b}$ $\frac{1}{2}b^{3}-\frac{1}{2}b^{2}$ Solve Upper Triangular System $\frac{1}{2}n_{b}^{2}-\frac{1}{2}n_{b}$ $\frac{1}{2}b^{3}-\frac{1}{2}b^{2}$ Full Matrix Multiplication $\frac{1}{3}n_{b}^{3}-\frac{1}{2}n_{b}^{2}+\frac{1}{6}n_{b}$ $b^{3}$ $\frac{1}{6}n_{b}^{3}-\frac{1}{2}n_{b}^{2}+\frac{1}{3}n_{b}$ $b^{3}$ Gramian Matrix Multiplication $\frac{1}{2}n_{b}^{2}-\frac{1}{2}n_{b}$ $\frac{1}{2}b^{3}$ Total Complexity $\frac{1}{3}m^{3}-\frac{1}{2}m^{2}+\frac{1}{6}m$ $\frac{1}{6}m^{3}-\frac{1}{4}m^{2}+\frac{1}{12}m$
The complexity of each operation is given by the number of times it is performed (a function of $n_{b}$) multiplied by the FLOP count per operation (a function of the block size $b$). Substituting $b=m/n_{b}$ into these expressions, the total complexity, obtained by summing across all operations, simplifies such that $n_{b}$ cancels out, as shown in the last row of Table E.1. Thus, the total computational complexity of MEMDET is independent of the number of blocks $n_{b}$, and is identical to that of conventional factorization algorithms where $n_{b}=1$.
Although the total computational cost remains the same, the contribution of individual operations shifts as $n_{b}$ increases. When $n_{b}=1$, the entire computation consists solely of a matrix decomposition. As $n_{b}$ increases, the decomposition cost decreases while additional operations, such as solving triangular systems and matrix multiplications, account for a larger fraction of the total complexity. At the extreme case of $n_{b}=m$, the algorithm consists primarily of matrix multiplications. This transition is illustrated in Figure E.1 (left panel), where the contributions of matrix decomposition, triangular system solving, and matrix multiplication are plotted as functions of $n_{b}$. The total computational complexity, shown as the black curve, remains constant, while the distribution of work among different operations shifts as $n_{b}$ increases.
Figure E.1: Theoretical computational complexity of MEMDET as a function of the number of blocks $n_{b}$. The left panel shows the contributions of matrix decomposition, solving triangular systems, and matrix multiplication to the total complexity. The black curve represents the total computational cost, which remains constant, while the colored curves illustrate how the workload shifts across operations as $n_{b}$ increases. The right panel displays the total data transfer volume (normalized by the original matrix size) for different $n_{b}$, highlighting the increasing cost of disk I/O as the number of blocks grows.
E.2 Data Transfer and Memory Considerations
While the total FLOP count is independent of $n_{b}$, the number of data transfers between memory and disk increases with the number of blocks. Table E.2 summarizes the number of blocks read from disk to memory and written back to disk, as a function of $n_{b}$. The actual volume of transferred data is obtained by multiplying the number of blocks by the block size, $b^{2}\beta$, where $\beta$ represents the number of bytes per floating point. The right panel of Figure E.1 illustrates the total data transfer volume relative to the original matrix size.
Table E.2: Number of blocks transferred between disk and memory during MEMDET execution. The total data transfer volume is obtained by multiplying the number of transferred blocks by the block size, $b^{2}\beta$ bytes, where $b=m/n_{b}$. Read operations occur in all cases, while write operations to the scratchpad are only required for $n_{b}>2$.
Operation Generic Matrix Symmetric Matrix
Read $\frac{2}{3}n_{b}^{3}-n_{b}^{2}+\frac{4}{3}n_{b}$ $\frac{1}{3}n_{b}^{3}-\frac{1}{2}n_{b}^{2}+\frac{7}{6}n_{b}$
Write $\begin{cases}0,&n_{b}\leq 2\\ \frac{1}{3}n_{b}^{3}-\frac{4}{3}n_{b}-1,&n_{b}>2\end{cases}$$\begin{cases}0,&n_{b}\leq 2\\ \frac{1}{6}n_{b}^{3}+\frac{1}{2}n_{b}^{2}-\frac{11}{3}n_{b}+4,&n_{b}>2\end{cases}$
For $n_{b}\leq 2$, the entire computation can often fit in memory, avoiding disk I/O, requiring no scratchpad space. However, for $n_{b}>2$, the data transfer overhead increases approximately as $\mathcal{O}(n_{b}^{2})$, meaning that choosing an excessively high $n_{b}$ introduces unnecessary I/O costs. Despite this, the hierarchical design of MEMDET efficiently schedules block transfers, mitigating excessive data movement.
Table E.3 provides an analysis of the required memory and scratchpad space. The number of blocks stored in memory and on disk is determined by $n_{b}$, with the actual space usage obtained by multiplying the number of blocks by the block size. By adjusting $n_{b}$, MEMDET can be configured to run within any given memory constraint, making it adaptable to systems with limited memory.
Table E.3: Number of concurrent blocks that must be allocated in memory and the total number of blocks allocated on disk during MEMDET execution. The total required memory and disk space are obtained by multiplying the number of allocated blocks by the block size, $b^{2}\beta$ bytes. While the number of concurrent memory-resident blocks remains fixed, the total number of blocks allocated on disk increases with $n_{b}>2$.
Hardware Generic Matrix Symmetric Matrix
Memory 3 or 4 3 or 4
Scratchpad $\begin{cases}0,&n_{b}\leq 2\\ n_{b}^{2}-n_{b}-1,&n_{b}>2\end{cases}$$\begin{cases}0,&n_{b}\leq 2\\ \frac{1}{2}n_{b}^{2}+\frac{1}{2}n_{b}-4,&n_{b}>2\end{cases}$
E.3 Empirical Performance Evaluation
To validate the theoretical complexity and memory analysis, we conducted empirical evaluations on SPD matrices of various sizes, ranging from $m=2^{10}$ to $2^{16}$. The largest matrix tested was chosen to match the memory capacity of a 64 GB system, allowing for a direct comparison between MEMDET and conventional algorithms. For each matrix size, the algorithm was executed with different numbers of blocks, $n_{b}=1,2,\dots,8$, where $n_{b}=1$ corresponds to a standard full-matrix decomposition with the entire matrix loaded into memory. Each experiment was repeated 10 times, and the mean and standard deviation of the profiling measures are reported.
Figure E.2 presents the experimental results. The left panel shows peak memory allocation, measured using a memory profiling tool, which precisely matches the theoretical predictions. As expected, when $n_{b}=1$, the required memory equals the original matrix size, while increasing $n_{b}$ reduces the memory footprint. The middle and right panels display the measured process time as functions of $n_{b}$ and $m$, respectively. At large $m$, the difference in process time across varying $n_{b}$ diminishes, indicating that the increase in data transfer cost does not significantly impact overall runtime.
Figure E.2: Peak memory allocation (a) and CPU processing time (b, c) for MEMDET on symmetric positive-definite matrices of size $m=2^{10},\dots,2^{16}$, using Algorithm 4. The matrices were processed using an $n_{b}\times n_{b}$ grid of matrix blocks, where $n_{b}=1,2,\dots,8$.
To further analyze this effect, Figure E.3 decomposes the process time into computation time and data transfer time. At small $m$, the runtime is dominated by disk I/O, but as $m$ increases, computation time becomes the dominant factor. This confirms that for sufficiently large matrices, the performance of MEMDET approaches that of conventional in-memory methods.
Figure E.3: Breakdown of MEMDET runtime (using Algorithm 4) into computation and data transfer times, normalized by $m^{3}$, for $m=2^{13},\dots,2^{16}$. The total process time consists of reading from disk to memory (maroon), writing from memory to disk (light tan), and computation (ochre).
E.4 Conclusing Remarks of Performance Analysis
MEMDET maintains the same computational complexity as conventional factorization methods while distributing computations across blocks. The total FLOP count remains unchanged, but increasing $n_{b}$ shifts the workload between operations (i.e., , from matrix decompositions to matrix multiplications). However, increasing $n_{b}$ also increases data transfer overhead, requiring a balance between reducing memory usage and minimizing disk I/O.
The hierarchical scheduling of block operations optimizes memory usage while limiting unnecessary data movement, ensuring that MEMDET remains efficient under constrained memory conditions. By allowing users to specify a memory limit, MEMDET enables the processing of arbitrarily large matrices on systems with any limited memory size.
For large-scale applications, where conventional methods exceed memory capacity, MEMDET provides a practical alternative. The experiments confirm that while data transfer overhead exists, it does not significantly impact runtime at large $m$, making MEMDET a viable solution for large-matrix computations on standard hardware.
Appendix F Background Material for Neural Scaling Laws
The noiseless result from (Li et al., , 2023) that we require is contingent upon the following assumptions:
Assumption F.1 (Eigenvalue Decay).
There exists a $\beta>1$ and constants $c_{\beta},C_{\beta}>0$ such that
$c_{\beta}i^{-\beta}\leq\lambda_{i}\leq C_{\beta}i^{-\beta},$ (F.1)
where the $\lambda_{i}$ are eigenvalues of the kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ under the decomposition guaranteed by Mercer’s Theorem:
$\kappa(x,x^{\prime})=\sum_{i=1}^{\infty}\lambda_{i}e_{i}(x)e_{i}(x^{\prime}).$ (F.2)
We need to define an embedding index associated to certain interpolation spaces that arise as the range of fractional powers of integral operators. In order to do so, we define the integral operator $T:L^{2}\to L^{2}$ that acts as the natural embedding of a RKHS $\mathcal{H}$ associated with our kernel $\kappa$, precomposed with its adjoint. That is, $T$ is the integral operator given by
[…]
where $\mu$ is the marginal distribution on $\mathcal{X}$ of $\rho$, the source distribution on $(\mathcal{X}\times\mathcal{Y})$ underlying the dataset. The operator $T$ can be decomposed by the spectral theorem of compact self-adjoint operators via
$T=\sum_{i=1}^{\infty}\lambda_{i}\langle\cdot,e_{i}\rangle_{L^{2}}e_{i}.$ (F.3)
For $s\geq 0$, this lets us define the fractional powers $T^{s}:L^{2}\to L^{2}$ of the operator $T$ to satisfy
$T^{s}(f)=\sum_{i=1}^{\infty}\lambda_{i}^{s}\langle f,e_{i}\rangle_{L^{2}}e_{i}.$ (F.4)
The interpolation space $[\mathcal{H}]^{s}$ associated to $T^{s/2}$ can then be defined as
$[\mathcal{H}]^{s}=\mathrm{Range}(T^{s/2})=\left\{\sum_{i=1}^{\infty}a_{i}\lambda_{i}^{\frac{s}{2}}e_{i}\mid\sum_{i=1}^{\infty}a_{i}^{2}<\infty\right\}\subset L^{2}.$ (F.5)
We now say that $\mathcal{H}$ has an embedding property of order $\alpha\in(0,1]$ if $[\mathcal{H}]^{\alpha}$ can be continuously embedded into $L^{\infty}$. Define then the operator norm, which has the form (see (Fischer & Steinwart, , 2020))
$\|[\mathcal{H}]^{s}\hookrightarrow L^{\infty}\|=\mathrm{ess}\sup_{x\in\mathcal{X},\mu}\sum_{i=1}^{\infty}\lambda^{\alpha}_{i}e_{i}(x)^{2}.$
We now have the following assumption on the embedding index, which is known to be satisfied if the eigenfunctions $e_{i}$ are uniformly bounded (Steinwart et al., , 2009).
Assumption F.2 (Embedding index).
The embedding index $\alpha_{0}=1/\beta$, where $\beta$ is the eigenvalue decay in (F.1), and $\alpha_{0}$ is defined as
$\alpha_{0}=\inf\left\{\alpha:\|[\mathcal{H}]^{s}\hookrightarrow L^{\infty}\|=M_{\alpha}<\infty\right\}$
Finally, we have the following assumption on the smoothness of the source function $f^{\star}_{\rho}=\mathbb{E}_{\rho}[y\mid x]$, which is a more precise characterization than requiring it to belong to some interpolation space.
Assumption F.3 (Source condition).
There exists an $s>0$ and a sequence $(a_{i})_{i\geq 1}$ for which
$f^{\star}_{\rho}=\sum_{i=1}^{\infty}a_{i}\lambda_{i}^{\frac{s}{2}}i^{-\frac{1}{2}}e_{i},$
and $0<c\leq|a_{i}|\leq C$ for some constants $c,C$.
These assumptions are required for the following theorem, taken from (Li et al., , 2023).
Theorem F.1.
Under Assumptions F.1, F.2 and F.3, fix $s>1$ and suppose that $\lambda\asymp n^{-\theta}$ for $\theta\geq\beta$. Then
$\mathbf{\mathrm{Bias}}^{2}=E(n)=\mathcal{O}^{\mathrm{poly}}_{\mathbb{P}}(n^{-\min(s,2)\beta}).$
Appendix G Proofs of Lemma 1 and Proposition 1
  • Proof of Lemma 1.
    Fix $\alpha>0$ and write $\mathbf{K}_{n}^{\alpha}=\mathbf{K}_{n}+\alpha\mathbf{Id}_{n\times n}$. In block form, $\mathbf{K}_{n}^{\alpha}$ contains $\mathbf{K}_{n-1}^{\alpha}$ according to
    $\mathbf{K}_{n}^{\alpha}=\begin{bmatrix}\mathbf{K}_{n-1}^{\alpha}&\kappa(\boldsymbol{x}_{n-1},x_{n})\\ \kappa(\boldsymbol{x}_{n-1},x_{n})^{\intercal}&\kappa(x_{n},x_{n})+\alpha\end{bmatrix},$
    where $\kappa(\boldsymbol{x}_{n-1},x_{n})=(\kappa(x_{i},x_{n}))_{i=1}^{n-1}\in\mathbb{R}^{nd\times d}$. Consequently, since $\mathbf{K}_{n}^{\alpha}$ and $\mathbf{K}_{n-1}^{\alpha}$ are both positive-definite, their determinants differ by the Schur determinant:
    $\det(\mathbf{K}_{n}^{\alpha})=\det(\mathbf{K}_{n-1}^{\alpha})\det\big{(}\kappa(x_{n},x_{n})+\alpha-\kappa(\boldsymbol{x}_{n-1},x_{n})^{\intercal}[\mathbf{K}_{n-1}^{\alpha}]^{-1}\kappa(\boldsymbol{x}_{n-1},x_{n})\big{)}.$
    Combining the Sylvester rank inequality with Corollary 20 from (Ameli & Shadden, , 2023), we take $\alpha\downarrow 0$ and observe that
    $\mathsf{pdet}(\mathbf{K}_{n})\leq\mathsf{pdet}(\mathbf{K}_{n-1})\;\mathsf{pdet}(\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)).$ (G.1)
    This lets us apply the AM-GM inequality and then bound the Frobenius norm in terms of the nuclear norm to obtain
    $\frac{\mathsf{pdet}(\mathbf{K}_{n})}{\mathsf{pdet}(\mathbf{K}_{n-1})}$ $\leq\mathsf{pdet}(\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1))=\prod_{j=1}^{r}\lambda_{j}\leq r^{-\frac{r}{2}}\left(\sum_{j=1}^{r}\lambda_{j}^{2}\right)^{\frac{r}{2}}$
    $\leq r^{-\frac{r}{2}}\mathsf{trace}\left(\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)^{r}\right)$
    $\leq\left(\frac{d}{r}\right)^{r/2}E(n)^{r},$
    where the $\lambda_{j}$ are the non-zero eigenvalues of the covariance matrix, and $r$ is its rank.
  • Proof of Proposition 1.
    From equation (3),
    $\mathsf{logdet}(\mathbf{K}_{n})-\mathsf{logdet}(\mathbf{K}_{n-1})$ $=\log C-\nu\log n+\log[1+o_{p}(1)]$
    $=\log C-\nu\log n+o_{p}(1),$ (G.2)
    and so
    $\mathsf{logdet}(\mathbf{K}_{n})-\mathsf{logdet}(\mathbf{K}_{1})=(n-1)\log C-\nu\log(n!)+o_{p}(n).$
    Letting $c_{0}\coloneqq\log C-\mathsf{logdet}(\mathbf{K}_{1})$ and dividing by $n$ implies the first result. For the second result, we replace the $o_{p}(1)$ term in (G.2) with $\delta_{n-1}$. Consequently,
    $\frac{n}{\sqrt{n-1}}\left[L_{n}-L_{1}-\left(1-\frac{1}{n}\right)c_{0}+\nu\frac{\log(n!)}{n}\right]=(n-1)^{-1/2}\sum_{i=1}^{n-1}\delta_{i},$
    and from (Billingsley, , 1961), $(n-1)^{-1/2}\sum_{i=1}^{n-1}\delta_{i}$ converges weakly to a normal random variable with zero mean and variance $\sigma^{2}=\mathbb{E}[\delta_{1}^{2}]$. ∎
Appendix H Derivation of FLODANCE Parameterization
Letting $\nu_{n}=\nu_{0}+\sum_{i=1}^{q}\nu_{i}n^{-i}$, Proposition 1 implies that, asymptotically in $n$,
[…]
for some $c_{0},\nu_{0},\dots,\nu_{q}$ and $\sigma>0$. Rearranging, there is
$\frac{n}{\sqrt{n-1}}(L_{n}-L_{1})=\frac{n-1}{\sqrt{n-1}}c_{0}-\left(\nu_{0}+\sum_{i=1}^{q}\frac{\nu_{i}}{n^{i}}\right)\frac{\log(n!)}{\sqrt{n-1}}+\epsilon_{n},$
and equivalently,
$y_{n}=c_{0}x_{n,0}+\sum_{i=1}^{q+1}\nu_{i-1}x_{n,i}+\epsilon_{n},$
where
$x_{n,0}=\sqrt{n-1},\qquad x_{n,i}=\frac{\log(n!)}{n^{i-1}\sqrt{n-1}},\qquad y_{n}=\frac{n}{\sqrt{n-1}}(L_{n}-L_{1}).$
To validate the theoretical parameterization, we empirically compute the log-determinants of NTK submatrices of size $1,\dots,n$ from ResNet9 and ResNet50 trained on CIFAR-10, where $n=$50000$$. The log-determinants are computed using MEMDET (Algorithm 3) with LDL decomposition. Figure H.1 presents the empirical results, where the black curves (overlaid by orange curves) represent the computed log-determinants, and the orange curves show the theoretical fits obtained using the parameterization in Algorithm 1. The fitting is performed globally over the entire interval, demonstrating the effectiveness of the theoretical model in capturing the log-determinant behavior. The near-perfect overlap between the curves in both cases highlights the accuracy of the fit, with errors remaining below 0.05% for most of the interval.
Figure H.1: Log-determinant $\ell_{n}$ for $n=1,\dots,$50000$$, corresponding to $m\times m$ NTK submatrices where $m=nd$ and $d=10$, from ResNet9 (a) and ResNet50 (b) trained on CIFAR-10 with $50000$ datapoints. Values are computed using MEMDET (Algorithm 3) with LDL decomposition (black curves, overlaid by colored curves). The orange curves represent theoretical fits based on the parametrization derived in Algorithm 1. Fitting is performed globally over the entire interval $(n_{0},n)=(1,5\times 10^{4})$, demonstrating the accuracy of the theoretical model. The blue curve, corresponding to the right axis (scaled to one-thousandth of the left axis), shows the absolute error.
Appendix I Implementation and Reproducibility Guide
We developed a Python package detkit444detkit is available for installation from PyPI (https://pypi.org/project/detkit), the documentation can be found at https://ameli.github.io/detkit, and the source code is available at https://github.com/ameli/detkit. that implements the MEMDET algorithm and can be used to reproduce the numerical results of this paper. A minimalistic usage of the detkit.memdet function is shown in LABEL:list:detkit1, where the user can specify various parameters: the maximum memory limit (max_mem), the structure of the matrix via the assume argument—set to gen for generic matrices (Algorithm 2), sym for symmetric matrices (Algorithm 3), and spd for symmetric positive-definite matrices (Algorithm 4)—whether the data is provided in full or in its lower/upper triangular form (triangle), the arithmetic precision used during computation (mixed_precision), the location of scratchpad space on disk (scratch_dir), and enabling parallel data transfer between memory and disk (parallel_io).
Listing I.1: A minimalistic usage of detkit package. The function memdet computes $\mathsf{logabsdet}(\mathbf{M})$ using the MEMDET algorithm.
[…]
from detkit import memdet
import zarr
[…]
6M = zarr.open(’filename.zarr’, ’r’)
[…]
ld, sign, diag, perm, info = memdet(M, max_mem=’32GB’, assume=’sym’, triangle=’u’,
11 overwrite=False, mixed_precision=’float64’,
scratch_dir=’/tmp’, parallel_io=’tensorstore’,
verbose=True, return_info=True, flops=True)
An example of utilizing FLODANCE is shown in LABEL:list:detkit2 using the FitLogdet class in detkit. A small subset of log-determinants from submatrices, computed in LABEL:list:detkit1, is used to extrapolate log-determinants of larger submatrices via Algorithm 1.
Listing I.2: The class FitLogdet fits and extrapolates log-determinants using FLODANCE in Algorithm 1.
1import numpy as np
from detkit import FitLogdet
# Range of datapoints ;#(\(n\))#; and number of labels ;#\((d)\)#; for CIFAR-10
n, d = np.range(50000), 10
6
[…]
ell = np.cumsum(np.log(np.abs(diag)))
11# Keep every ;#\$d\$#;-th element
ell = ell[(d-1)::d]
# Choose a fit interval, such as ;#\$(n_0, n_s) = (10^2, 5 $\times$ 10^3)\$#;
n0, ns = 1e2, 5e3
16fit_mask = (n > n0) & (n < ns)
[…]
flodet = FitLogdet(q=4)
flodet.fit(n[fit_mask], ell[fit_mask])
21
# Extrapolate in a larger interval, such as in ;#\$ (n_s, n) = (5 $\times$ 10^3, 5 $\times$ 10^4)\$#;
n_eval = np.geomspace(ns, n)
ell_eval = flodet.eval(n_eval)
We have also concurrently developed a separate high-performance Python package, imate,555imate is available for installation from PyPI (https://pypi.org/project/imate), the documentation can be found at https://ameli.github.io/imate, and the source code is available at https://github.com/ameli/imate. which implements stochastic Lanczos quadrature (SLQ), a randomized method for approximating the log-determinant at scale. This package is implemented with a C++/CUDA backend and supports execution on both CPU and multiple GPUs. LABEL:list:imate demonstrates a minimalistic usage of the imate.logdet function.
Listing I.3: A minimalistic usage of imate package. The function logdet computes $\mathsf{logdet}(\mathbf{M})$ using the stochastic Lanczos quadrature algorithm.
[…]
from imate import logdet
import numpy
# Number of detests (;#\$ n \$#;) and labels (;#\$ d \$#;)
6n, d = 50000, 10
[…]
M = numpy.memmap(’filename.npy’, mode=’r’, dtype=’float32’, shape=(n*d, n*d))
[…]
ld, info = logdet(M, method=’slq’, min_num_samples=100, max_num_samples=200,
lanczos_degree=100, error_rtol=0.01, confidence_level=0.95,
outlier_significance_level=0.001, orthogonalize=-1, num_threads=0,
16 num_gpu_devices=0, gpu=True, verbose=True, return_info=True,
plot=True)